### v1

# Mix-of-Granularity: Optimize the Chunking Granularity for Retrieval-Augmented Generation

Abstract
Integrating information from different reference data sources is a major challenge for Retrieval-Augmented Generation (RAG) systems because each knowledge source adopts a unique data structure and follows different conventions. Retrieving from multiple knowledge sources with one fixed strategy usually leads to under-exploitation of information. To mitigate this drawback, inspired by Mix-of-Expert, we introduce Mix-of-Granularity (MoG), a method that dynamically determines the optimal granularity of a knowledge database based on input queries using a router. The router is efficiently trained with a newly proposed loss function employing soft labels. We further extend MoG to Mix-of-Granularity-Graph (MoGG), where reference documents are pre-processed into graphs, enabling the retrieval of relevant information from distantly situated chunks. Extensive experiments demonstrate that both MoG and MoGG effectively predict optimal granularity levels, significantly enhancing the performance of the RAG system in downstream tasks. The code of both MoG and MoGG will be made public.
1 Introduction
Retrieval-Augmented Generation (RAG) [30] has become a popular method for enhancing Large Language Models (LLMs). The core concept of RAG involves retrieving relevant information from external knowledge bases to provide additional context to the LLM, enabling it to generate more precise and grounded responses. RAG offers a promising and practical solution to mitigate LLMs’ hallucinations because (1) it can be applied to any LLM, even those accessible only via APIs, and (2) the reference information is easy to modify or update. Many LLM-based products are supported by RAG systems, with examples spanning various industries such as customer service, advertising and marketing, education and e-learning, healthcare, and e-commerce and retailing [27, 51, 44, 26, 36].
The quality and pertinence of the retrieved reference documents are crucial for the quality of the final generated response. Consequently, much research has focused on the retrieval phase. Currently, most RAG systems follow the Dual-Encoder Architecture [11] (DEA) paradigm, in which the reference documents are divided into small snippets, encoded by specific encoders, and then stored in the vector database (e.g. FAISS [24] or Neo4j [8]) as embeddings. The DEA paradigm shows great potential for empowering LLMs with any existing knowledge database (knowledge graphs, textbooks, Wikipedia, etc.), providing a practical solution to alleviate LLMs’ hallucinations.
However, a notable challenge in adopting the DEA paradigm is that the chunking size of reference documents is hard to optimize. For instance, effectively retrieving reference knowledge from medical textbooks or a Medical Knowledge Graph Database (e.g. Hetionet [16]) requires different chunking sizes because the data structure and information density in both sources are quite different: medical textbooks contain long text passages organized around similar topics, while a Medical Knowledge Graph Database is formed with entities of different terms (usually quite short) and the relations between them. In practice, different chunking sizes should be set for these two knowledge databases. However, searching for the optimal chunking size for each knowledge source is tedious. Another challenge of the DEA paradigm is that even when an optimal chunking size is determined, this fixed chunking size does not work well in all cases, because different types of users’ input correspond to different optimal chunking sizes. As illustrated in Figure 1, when the user asks for precise information about one disease or one medicine (fine-grained question), chunking the reference document in finer granularity is better; whereas, if the user asks for broad information, like the general information of a disease (coarse-grained question), a more coarse granularity of chunking is preferred. Thus, the community calls for a method to dynamically determine the chunking size, for which we propose the Mix-of-Granularity (MoG). We draw inspiration from Mix-of-Experts [6], which is a machine learning architecture that dynamically combines predictions from multiple specialized models, allowing for adaptive selection of the most relevant expertise for each input. Similarly, the fundamental concept of MoG involves training, in a supervised learning manner, a router to select and combine reference snippets of different granularity levels so that only the most pertinent snippets are passed to the backbone LLM. With the snippets of improved quality, the LLM can give better-grounded responses.
Figure 1: The Mix-of-Granularity (MoG) mechanism can automatically select the optimal granularity when extracting information from the reference knowledge database (scenarios 1 and 2), achieving both high information pertinence and coverage. The selected text is passed to LLM for generation. When relevant information is dispersed across distant sections (scenario 3), MoG-Graph (MoGG) is applied to effectively retrieve these separate pieces of information. In MoGG, the reference documents are pre-processed in the form of a graph, and MoGG retrieves information with the best hopping range. The text corresponding to the selected subgraph is then fed to LLM for RAG.
The MoG still has difficulty dealing with broader queries that require information from different documents of a single knowledge database or even information from different knowledge databases. In such cases, i.e., cross-document question in Figure 1, adjusting the granularity of the reference document is not helpful because the necessary information is so distantly located that it cannot be covered by a reasonably large chunking window. To better answer these broader queries, we extend MoG to Mix-of-Granularity-Graph (MoGG). In MoGG, the reference documents in the knowledge databases are pre-processed as a graph, allowing relevant snippets to be included as neighbors of each other, regardless of their distance in the original databases. This extension further improves the performance of retrieving information from multiple knowledge sources.
In most existing RAG systems, when retrieving information from a database, a top-$k$ selection is implemented to control the number of candidate snippets. This common practice obstructs the backward propagation of the gradient, making the training infeasible. In this paper, we keep this top-$k$ selection for maximum compatibility with other works. To solve the problem of backward propagation, we introduce a loss function using soft labels. Soft labels are approximate training signals generated using offline algorithms or models like TF-IDF [38] or RoBERTa [32]. With the soft labels, the top-$k$ retrieval is excluded from the training process, thus the issue of backward propagation is circumvented and the training is accelerated.
In conclusion, the main contributions of this work are: (1) We propose MoG, which dynamically retrieves snippets from the optimal granularity level with the help of a router, achieving a balanced trade-off between coverage and pertinence during the retrieval phase of any RAG system. (2) We extend MoG to MoGG by reorganizing the reference document in the form of a graph, thereby further improving the quality of retrieval when relevant information is dispersed across distant sections in the knowledge database. (3) We introduce a loss function utilizing soft labels to overcome the challenges associated with training with top-$k$ selection.
2 Related Work
2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) [30] has emerged as a standard practice to enhance the performance of LLMs, aiming to mitigate their problems of “hallucinations” and knowledge cut-off. An RAG system typically includes a Retriever that extracts relevant information from an external knowledge database, and a backbone LLM to generate grounded responses by considering the given relevant information (in-context learning [10]).
Technological advancements in RAG have addressed fundamental questions such as “what should be retrieved”, “when should retrieval occur”, and “how should the retrieved information be utilized”. In terms of “what to retrieve”, research has evolved from simple token retrieval [28] and entity retrieval [34] to more complex structures like chunks [37] and knowledge graphs [25]. Granularity matters a lot in retrieval, coarse-granularity-retrieval yields more information but with lower precision, while fine-granularity-retrieval offers comprehensive information at the cost of efficiency. The question of “when to retrieve” has led to strategies ranging from single [49, 42] to adaptive [20, 17] and multiple retrieval [19] methods. Regarding “how to use” the retrieved data, integration techniques have been developed for various levels of the model architecture, including the input [29], intermediate [3], and output layers [31]. Although integration at the intermediate and output layers is more effective, challenges remain concerning training requirements and efficiency limitations.
2.2 Chunking Optimization
Optimal chunk size is crucial for managing external documents. Breaking them down into small chunks is the first step to extracting and encoding the fine-grained features. Naive chunking strategies, such as “fixed-size” chunking and “recursive chunking,” attempt to create snippets of identical or similar size. However, fixed chunking size often yields suboptimal retrieval results in practice.
Current research in RAG explores chunking optimization techniques to improve retrieval efficiency and accuracy. One line of work focuses on increasing the recall of relevant snippets. For example, the “Sliding Window Chunking” [40] allows layered retrieval by merging globally related information across multiple processes. The “Parent Document Retrieval” [46] retrieves with small chunks and returns larger blocks of context for language model processing. Another line of work seeks to include more semantic information of the context to improve retrieval accuracy. “Metadata Filtering” [43] leverages document metadata to enhance the filtering process; “Context-Enriched Chunking” [45] breaks down information into meaningful segments and adds helpful summaries; while “Windowed Summarization Chunking” [45] enriches each text chunk with a windowed summary of the previous few chunks. Combining these methods has led to notable advancements, enhancing retrieval outcomes and improving performance in RAG.
2.3 Graph-Based Text Processing
Graph-based text processing techniques combine research in graphs and text retrieval. Previous works exploit semantic similarities between small snippets (a sentence or several words) and reorganize the text material into a graph using Entity Recognition and Relation Construction algorithms [33, 14, 52]. This approach breaks the constraint of the single dimension in the original text corpus, allowing chunks of the same topic to be grouped as neighbors in a graph. These graph-based text processing methods show great potential in tasks requiring reasoning over long texts or multi-hop reasoning.
Representative works in this field include “abstract embedding” [13], which prioritizes top-$k$ retrieval based on document abstracts (or summaries), providing a comprehensive understanding of the context. RAPTOR [41] organizes snippets as a tree (a special form of a graph) by recursively clustering them, where all non-leaf nodes correspond to summaries of their child nodes. This processing allows access to information at different granularity levels, resulting in significant improvements in summarization tasks. In GMoE [48], the authors use different expert networks to handle hop-$1$, hop-$2$, and mixed hop-$1$ hop-$2$ neighbors of a node in a graph, which inspires our extension from MoG to MoGG.
3 Methodology
3.1 Preliminaries
Before describing the details of MoG and MoGG, let us begin with the basic annotations. MoG and MoGG are designed to enhance an existing RAG system, particularly its Retriever module. A typical RAG system comprises a Retriever $\mathcal{R}$, a Generator $\mathcal{G}$, and an external Knowledge database $\mathcal{K}$ (knowledge source). A RAG system takes a user query $q$ as input, retrieves reference document pieces (snippets or chunks) from $\mathcal{K}$, and uses these snippets to help the $\mathcal{G}$ produce the final responses.
In practice, the most popular architecture for the Retriever is the Dual-Encoders Architecture [11], where the query $q$ and all the snippets in $\mathcal{K}$ are encoded into embeddings ($e_{q}$ and $e_{s}$) using the same encoder $\mathcal{E}$. The extraction of relevant snippets is achieved by calculating the similarity between $e_{q}$ and $e_{s}$. The calculation of cosine similarities is made efficient thanks to modern vector databases like FAISS [24] or the Neo4j Database [8]. The Generator module is typically a backbone LLM, such as GPT [4], Llama [1], GLM [12], or InternLM [5]. Once the snippets are retrieved, they are usually injected into the backbone LLM via prompt injection. This method is favored because many powerful LLMs are accessible only via API calls. With the knowledge-integrated prompt, the LLM generates the final response. In some RAG systems, post-processing modules are added after generation, but this aspect is not covered in our work as it is not our focus.
3.2 Naive MoG
3.2.1 Multi-granularity Router
We apply the idea of Mix-of-Expert [6] (MoE) to automatically determine the best granularity level in the retrieval phase of RAG. In a MoE system, different input tokens will be routed to the best expert network to be processed efficiently. Similarly, in MoG, a router is trained to predict the importance weight of different granularity levels based on the user’s input, so that the information retrieved from the optimal granularity level can be prioritized in the final top-$k$ selection and thus has a higher chance to be passed to the LLM.
Figure 2 illustrates the workflow of retrieving relevant snippets with a router. At the beginning, the user’s input query $q$ is encoded via RoBERTa before being mapped to a weight $w$ by the router. In parallel, the input query is used to retrieve relevant snippets from the knowledge database $\mathcal{K}$. In each $\mathcal{K}$, the documents within it are first chunked into $n_{gra}$ different granularity levels. Each snippet is then encoded into embeddings. The same retriever will retrieve the $k_{r}$ most relevant snippets (in Figure 2 $k_{r}$=3) from each granularity level based on their similarity with $q$. After the retrieval, a pool of candidate snippets of size $n_{gra}\times k_{r}$ is formed. During the retrieval, each candidate snippet is assigned a relevance score. Since all snippets are retrieved with the same retriever, their relevance scores are comparable. These scores are then weighted and summed using weight $w$ generated by the router. The $k$ snippets with the highest weighted relevance scores ($k$=1 in Figure 2) are selected and fed to the backbone LLM, helping LLM generate grounded responses.
Figure 2: Mix-of-Granularity (MoG) mechanism prioritizes the chunks retrieved from optimal granularity level, which is determined by the router based on the user input query.
To trace the relevance scores of the same snippet across the different granularity levels, the documents are chunked in a non-overlapped way, and all chunks of coarser granularity are concatenations of several chunks of the finer granularity. In other words, each snippet of the finest granularity is included in one and only one snippet of coarser granularity. This approach allows us to trace the relevance scores of different granularity levels by the IDs of the finest chunks. For each chunk of the finest granularity level, if it is retrieved among the top-$k_{r}$ snippets from one granularity level, its relevance score for that level is recorded as the actual relevance score, just like the chunk marked in red in Figure 2; otherwise, it is filled with 0, like some zeroes padded for the chunk marked in blue. In this way, we get the relevance scores $t_{rs}$ of each finest snippet. Noting the set of all chunks of the finest granularity level as $\mathcal{C}$, the retrieved chunk ($chunk_{r}$) is selected by top-$k$ selection as follows:
$chunk_{r}=\underset{c\in\mathcal{C}}{\operatorname{top_{k}-argmax}}(t_{rs}(c)\cdot w).$ (1)
The snippet injected to the prompt of LLM is the snippet containing $chunk_{r}$ of the granularity level with the highest weight ($g_{r}$). In Figure 2, the snippet of the third granularity level is injected in the prompt because this level has the highest weight. Formally, $g_{r}$ is determined by:
$g_{r}=\underset{g\in[1,n_{gra}]|w_{g}\neq 0}{\operatorname{argmax}}w_{g},$ (2)
where $w_{g}$ represents the $g$-th element in $w$.
3.2.2 Soft-labels
In MoG, we use a Multi-Layer Perceptron (MLP) as the router, which is trained in a supervised learning method. The input for the router is the embedding of the input query $q$ generated with RoBERTa [32], which is then mapped to $w$ by the router. The intrinsic training signal is the “labels” ($l$) in the Medical Question-Answering (MQA) datasets. In some MQA datasets, the ground truth snippets to be retrieved are provided, so we use them as “labels” directly; otherwise, the concatenation of the strings of the “Question” and “Answer” is used as the “label”. A natural training objective is to maximize the semantic similarity between $chunk_{r}$ and $l$ by adjusting $w$. Unfortunately, this label can not be used to guide the training directly because there is a non-differentiable top-$k$ selection in the way. The soft labels are proposed to bypass the top-$k$ selection during the training.
For each query $q$, the most relevant snippet ($S_{best}$) is retrieved from the reference documents of each granularity level. The semantic similarity between each snippet in $S_{best}$ and the label $l$ is then calculated (with static models including TF-IDF [38], RoBERTa [32], or hitrate score) and stored in $sim_{best}$. We create a soft label of 0.8 (resp. 0.2) for the most (resp. the second) similar snippet in $S_{best}$, and pad 0 for the other snippets.
With the soft labels ($sl$) built, we can train the router to predict a high value (0.8) for the optimal granularity level, while conserving certain flexibility to choose the second-best granularity level. The router is trained by minimizing a Binary Cross Entropy loss function ($l_{bce}$):
$l_{bce}=\sum_{i\in len(w)}-[sl_{i}\cdot\log(w_{i})+(1-sl_{i})\cdot\log(1-w_{i})].$ (3)
3.3 MoGG: MoG with Graph-context
In MoG, by adjusting the granularity level, the adjacent snippets with relevant knowledge can be retrieved altogether. This method is particularly effective when information centered around the same topic is stored in sentences close to each other. However, in most cases, answering a complex question requires reasoning over information stored in different paragraphs or even different documents. In practice, the solution is to perform more retrievals at a finer granularity level, retrieving only highly relevant small pieces to form a comprehensive reasoning chain. This approach is inconvenient because determining the optimal number of retrievals often involves manually adjusting $k$ for the final selection, which is challenging because the whole tuning process is time-consuming and $k$ can not go infinitely large. In this scenario, MoG shows no clear advantage compared to classic RAG implementations, because when changing the $k$ is hard, adjusting the granularity level will only sacrifice either the precision or recall of the retrieval.
To overcome this challenge, an intuitive approach is to reorganize the reference document and group the relevant information so that each snippet in Figure 2 contains both the originally dispersed relevant text and the adjacent content. In this way, even for a limited $k$, the selected snippets can encompass information from different paragraphs or documents. Motivated by this, we propose a more applicable framework, MoGG, by enhancing MoG with a preprocessing step that organizes the reference database $\mathcal{K}$ into a graph. As illustrated in Figure 3, each document is initially split into small pieces consisting of one or two sentences, and each piece is treated as a separate node in the graph. To determine the edges in the graph, an index is first created with all these nodes. Then, each node is used as a “query” to search for the $k_{graph}$ (set as 3 in Figure 3) most relevant nodes using a retriever. An edge is then added between the query node and each of the relevant nodes while ensuring that the similarity between them meets a predefined threshold $T_{graph}$. For simplicity, here we use the same retriever as the RAG system. The graph obtained is a binary undirected graph, meaning that the connections between any nodes are bi-directional.
Figure 3: Pre-processing the reference document to form graphs. The concept of “granularity level” is changed into “hopping range” in graphs.
This transformation of the original reference document into a graph extends the concept of “context” in linear text to “neighbors” in a graph. In a graph setting, granularity levels are adapted to “hopping ranges”: in a non-graph setting, a larger granularity level corresponds to more adjacent sentences being grouped in a chunk; in a graph setting, a larger granularity level corresponds to sentences within a larger hopping range of a centered sentence being grouped as one chunk. To avoid context redundancy, duplicate nodes in the neighbors are considered only once. Similar to MoG, the documents in the external knowledge database $\mathcal{K}$ are chunked with $n_{gra}$ different hopping ranges and then encoded into embeddings. The rest of the MoG, i.e., parts I and III in Figure 2, remains unchanged. By transforming the documents into a graph structure and defining granularity based on hopping ranges, MoGG effectively captures dispersed relevant information, allowing for more comprehensive and efficient retrieval.
4 Experiments
4.1 Corpus and Medical QA datasets
A reference knowledge database used in a RAG system is often termed a “corpus”. To form the corpora, data from various sources were collected, including the widely-used PubMed  [^1] https://pubmed.ncbi.nlm.nih.gov/ corpus for all biomedical abstracts, the StatPearls  [^2] https://www.statpearls.com/ corpus for clinical decision support, the Textbooks[21] corpus covering medical textbook knowledge, and the Wikipedia  [^3] https://www.wikipedia.org corpus for general knowledge. The four corpora are combined to form a larger corpus, named “MedCorp”. The statistics of the used corpora are shown in Table 1 (left).
Following the setup in the MIRAGE benchmark [51], we evaluate the performance of the RAG system using five Medical Question-Answering datasets: MMLU-Med [15], MedQA-US [53], MedMCQA [35], PubMedQA* [22], and BioASQ-Y/N [47]. Specifically, we extracted only the multiple-choice questions relevant to the biomedical field and removed the ground truth supporting contexts during testing. The details of each QA dataset are included in Appendix A. The statistics of the five datasets used in our experiments are listed in Table 1 (right).
Table 1: Important statistics of corpora and QA datasets
Corpus #Doc #Snip. $\bar{L}$. Domain
PubMed 23.9M 23.9M 196 Bio.-Med.
StatPearls 9.3k 301.2k 119 Clinics
Textbooks 18 125.8k 182 Medicine
Wikipedia 6.5M 29.9M 162 General
MedCorp 30.4M 54.2M 221 Mixed
Datasets Size #Opt. $\bar{L}$ Source
MMLU-Med 1,089 4 63 Exam.
MedQA-US 1,273 4 177 Exam.
MedMCQA 4,183 4 26 Exam.
PubMedQA* 500 3 24 Literature
BioASQ-Y/N 618 2 17 Literature
4.2 Experiment Setup
All backbone LLMs, whether accessed via API or local deployment, are run under off-the-shelf settings. The exact versions of the backbone LLMs are listed in Appendix B. Experiments are conducted on Nvidia GeForce 3090 and 4090 GPUs. The code is written using the PyTorch framework, utilizing an Adam optimizer with a learning rate of 0.001. Each training job is run until the convergence of the loss value. During the experiment, there are two top-k selections. Unless specified, when retrieving snippets from each corpus, we select the top-3 snippets; when all the snippets are retrieved from each corpus, we select the top-2 snippets with the highest relevance scores to pass to the backbone LLM. When snippets are too long, they are truncated automatically to fit the LLM’s context window size. The router requires approximately 12GB of GPU memory for training and 6GB for inference. Utilizing a caching mechanism, we efficiently completed 35 training sessions and over a hundred inferences, each training session taking around 4 hours for 1000 epochs.
4.3 Performance of MoG on Medical QA Task
As mentioned above, we test the effectiveness of MoG on Medical Question-Answering datasets. Each question within the dataset is formulated as a multiple-choice question. The LLM is tasked with choosing the best answer option(s) from the given options. Under this setting, we evaluated the performance of the entire RAG system by calculating the Exact Matching accuracy of the answers. To prevent knowledge leakage, following previous work [51], only the question is used (options not given) to retrieve reference documents from the external knowledge database.
The router of MoG guides the retrieval system to choose the optimal granularity. As shown in Figure 4, when tested on different datasets, the router trained with Textbook corpus shows a preference for different granularity levels. For instance, on the PubMedQA dataset, the finest granularity snippets are selected most frequently. This is because the questions in PubMedQA are typically precise and can be answered with short reference snippets.
Figure 4: Averaged weights of different granularity levels on different QA datasets
MoG is integrated into one same RAG system, with which we conduct the Medical Question-Answering task. Backbone LLMs are altered to cover some of the popular ones, such as ChatGPT [4], InternLM2 [5], Llama3 [1], GLM3 [12], and Qwen1.5 [2]. The router is trained with the soft labels built using RoBERTa [32], this choice is justified by the experiment in Appendix C. Three candidate snippets are retrieved from each granularity level of the external knowledge database. To investigate the effect of varying the number of candidate snippets on RAG system performance, we conducted experiments with different snippet counts. Detailed methodology and results are provided in Appendix D. The retriever is fixed as BM25 [39], with a further discussion on the performance of different retrievers included in Appendix E. Here we present the experimental results obtained with the router trained with MedCorp corpus in Table 2, the results obtained with routers trained on four single corpora are presented in Appendix F. The results demonstrate that MoG consistently enhanced the performance of the RAG system across different backbone models. Apart from the standard baseline using Chain-of-Thought (CoT) [50], we add a RAG system with a single granularity corpus as the baseline. The granularity level selected for this baseline is set to be the same as the one in MedRAG [51], therefore, it is also noted as “MedRAG” in Table 2.
From the table, we find that MoG constantly performs better than MedRAG, while not necessarily better than CoT. The reason why CoT might be better is that the RAG system we used has no noise filters, thus the noise is injected together with the external knowledge via prompts. A detailed analysis of the number of samples improved or degraded by the application of MoG is included in the Appendix G, in which we manually verified that the majority of degradation is caused by noise.
We also find that MoG improves the accuracy score more when applied on smaller, weaker LLMs (like ChatGLM and Qwen), this is probably because smaller LLMs have less knowledge stored in their internal parameters and, thus could benefit more from the retrieved snippets.
Table 2: Accuracy of Medical Question-Answering task with MoG (trained with MedCorp), best results marked in bold.
LLM Method MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
GLM3 CoT 0.4356$\pm$0.04 0.3451$\pm$0.03 0.3250$\pm$0.02 0.3400$\pm$0.05 0.5081$\pm$0.04 0.3908
MedRAG 0.4950$\pm$0.04 0.3804$\pm$0.03 0.3632$\pm$0.02 0.5100$\pm$0.05 0.6532$\pm$0.04 0.4804
MoG 0.5198$\pm$0.04 0.3529$\pm$0.03 0.3716$\pm$0.02 0.5400$\pm$0.05 0.6774$\pm$0.04 0.4923
GPT-3.5 CoT 0.7525$\pm$0.03 0.6118$\pm$0.03 0.5890$\pm$0.02 0.5200$\pm$0.05 0.7339$\pm$0.04 0.6571
MedRAG 0.6931$\pm$0.03 0.6510$\pm$0.03 0.4671$\pm$0.02 0.6000$\pm$0.05 0.8306$\pm$0.03 0.6484
MoG 0.7129$\pm$0.03 0.6471$\pm$0.03 0.5532$\pm$0.02 0.6200$\pm$0.05 0.7823$\pm$0.04 0.6631
InternLM CoT 0.7426$\pm$0.03 0.6118$\pm$0.03 0.5269$\pm$0.02 0.3500$\pm$0.05 0.7258$\pm$0.04 0.5914
MedRAG 0.6040$\pm$0.04 0.5294$\pm$0.03 0.3847$\pm$0.03 0.4300$\pm$0.05 0.7661$\pm$0.04 0.5428
MoG 0.7129$\pm$0.03 0.5961$\pm$0.03 0.5436$\pm$0.02 0.4100$\pm$0.05 0.7661$\pm$0.04 0.6057
Llama3 CoT 0.7079$\pm$0.03 0.6431$\pm$0.03 0.5663$\pm$0.02 0.5500$\pm$0.05 0.7258$\pm$0.04 0.6386
MedRAG 0.6040$\pm$0.03 0.5725$\pm$0.03 0.4313$\pm$0.02 0.5600$\pm$0.05 0.7823$\pm$0.04 0.5900
MoG 0.7228$\pm$0.03 0.6000$\pm$0.03 0.5627$\pm$0.02 0.6400$\pm$0.05 0.7984$\pm$0.04 0.6648
Qwen1.5 CoT 0.4604$\pm$0.04 0.3255$\pm$0.03 0.3883$\pm$0.02 0.2000$\pm$0.03 0.5484$\pm$0.04 0.3845
MedRAG 0.5594$\pm$0.03 0.4353$\pm$0.03 0.4038$\pm$0.02 0.3400$\pm$0.05 0.5403$\pm$0.04 0.4558
MoG 0.5941$\pm$0.04 0.4235$\pm$0.03 0.4301$\pm$0.02 0.4700$\pm$0.05 0.6694$\pm$0.04 0.5174
4.4 Performance of MoGG on Medical QA Task
Similarly to the experiment of MoG, we test the performance of MoGG on Medical QA datasets. Due to the large volume of Wikipedia and PubMed corpus, building the index for them in the setting of MoGG is too time-consuming. Therefore, we tested only the performance of MoGG with routers trained on Textbooks and StatPearls corpora. In Table 3 we present the results obtained when trained with Textbook corpus. The results obtained with StatPearls show similar patterns (listed in Appendix H). From the table, we can tell that MoGG can further improve the averaged accuracy scores. By comparing Table 2 and Table 3, we can find that, even when trained with significantly fewer samples (Textbooks corpus is a tiny subset of MedCorp corpus, accounting for only about 0.2% of all the snippets in MedCorp), MoGG brings more significant improvement in terms of the averaged accuracy score with respect to MedRAG than MoG. This finding highlights that MoGG is more efficient than MoG thanks to its flexible way of organizing the reference snippets (in the form of a graph).
Table 3: Accuracy of Medical Question-Answering task with MoGG (trained with Textbooks), best results marked in bold.
LLM Method MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
GLM3 CoT 0.4356$\pm$0.04 0.3451$\pm$0.03 0.3250$\pm$0.02 0.3400$\pm$0.05 0.5081$\pm$0.04 0.3908
MedRAG 0.4802$\pm$0.04 0.3569$\pm$0.03 0.3811$\pm$0.02 0.3600$\pm$0.05 0.5565$\pm$0.04 0.4269
MoG 0.5545$\pm$0.04 0.2941$\pm$0.03 0.3548$\pm$0.02 0.4700$\pm$0.05 0.5726$\pm$0.04 0.4492
MoGG 0.5347$\pm$0.04 0.3176$\pm$0.03 0.3608$\pm$0.02 0.4500$\pm$0.05 0.5645$\pm$0.04 0.4455
GPT-3.5 CoT 0.7525$\pm$0.03 0.6118$\pm$0.03 0.5890$\pm$0.02 0.5200$\pm$0.05 0.7339$\pm$0.04 0.6571
MedRAG 0.7277$\pm$0.03 0.6745$\pm$0.03 0.4468$\pm$0.02 0.2600$\pm$0.04 0.5161$\pm$0.04 0.5250
MoG 0.7525$\pm$0.03 0.6667$\pm$0.03 0.5603$\pm$0.02 0.5200$\pm$0.05 0.7016$\pm$0.04 0.6122
MoGG 0.7673$\pm$0.03 0.6784$\pm$0.03 0.5233$\pm$0.02 0.4700$\pm$0.05 0.6452$\pm$0.04 0.6168
InternLM CoT 0.7426$\pm$0.03 0.6118$\pm$0.03 0.5269$\pm$0.02 0.3500$\pm$0.05 0.7258$\pm$0.04 0.5914
MedRAG 0.6188$\pm$0.03 0.5569$\pm$0.03 0.3118$\pm$0.02 0.1400$\pm$0.03 0.4839$\pm$0.04 0.4223
MoG 0.7277$\pm$0.03 0.5725$\pm$0.03 0.5281$\pm$0.02 0.3400$\pm$0.05 0.7339$\pm$0.04 0.5804
MoGG 0.7228$\pm$0.03 0.5882$\pm$0.03 0.5173$\pm$0.02 0.3400$\pm$0.05 0.7339$\pm$0.04 0.5804
Llama3 CoT 0.7079$\pm$0.03 0.6431$\pm$0.03 0.5663$\pm$0.02 0.5500$\pm$0.05 0.7258$\pm$0.04 0.6386
MedRAG 0.6485$\pm$0.03 0.5961$\pm$0.03 0.4146$\pm$0.02 0.3800$\pm$0.05 0.5242$\pm$0.04 0.5127
MoG 0.7228$\pm$0.03 0.6196$\pm$0.03 0.5484$\pm$0.02 0.5100$\pm$0.05 0.7097$\pm$0.04 0.6221
MoGG 0.7030$\pm$0.03 0.5961$\pm$0.03 0.5460$\pm$0.02 0.5200$\pm$0.05 0.7661$\pm$0.04 0.6262
Qwen1.5 CoT 0.4604$\pm$0.04 0.3255$\pm$0.03 0.3883$\pm$0.02 0.2000$\pm$0.03 0.5484$\pm$0.04 0.3845
MedRAG 0.5941$\pm$0.03 0.4000$\pm$0.03 0.3835$\pm$0.02 0.3300$\pm$0.05 0.4919$\pm$0.04 0.4399
MoG 0.5792$\pm$0.03 0.3843$\pm$0.03 0.4110$\pm$0.02 0.3300$\pm$0.05 0.6129$\pm$0.04 0.4635
MoGG 0.5594$\pm$0.03 0.4314$\pm$0.03 0.4480$\pm$0.02 0.3000$\pm$0.05 0.6371$\pm$0.04 0.4752
5 Limitations and Broader Impacts
Our work serves as an early trial in dynamic chunking strategy, with the following major directions for improvement. First, MoG(G)’s candidate granularity levels are manually assigned. It could be more efficient if an algorithm automatically set these granularity levels to avoid excessive grid-searching for parameter optimization. Secondly, the current router uses only the semantic information of the input query to predict the best granularity level. In practice, other factors might influence this choice, such as the user’s knowledge of relevant fields, the type of query, and the expected response length. Incorporating more information into the router could potentially improve prediction results. Lastly, MoG(G) could be combined with other techniques to further enhance retrieval quality, such as Recursive Character Splitting  [^4] https://python.langchain.com/v0.1/docs/modules/data_connection/document_transformers/recursive_text_splitter, Parent Document Retrieval [46], or Sliding Window Chunking [40].
MoG(G) provides a valuable tool for integrating multiple sources of external knowledge databases into a RAG system, thereby enhancing RAG systems with existing knowledge sources. However, the router also introduces a new security risk: a compromised router could redirect knowledge retrieval to malicious sources, injecting incorrect or even harmful information into the backbone LLM. Therefore, it is crucial to protect and monitor the router to mitigate this risk.
6 Conclusion
In this work, we present Mix-of-Granularity (MoG), a mechanism to dynamically choose the best granularity when retrieving information from an external knowledge database. When applied to a Retrieval-Augmented Generation system, MoG helps retrieve more relevant information while reducing noise. MoG is further extended as Mix-of-Granularity-Graph (MoGG), where reference documents are reorganized in graph form. This extension allows distantly situated information to be retrieved simultaneously, overcoming the limitations of a fixed top-$k$ selection strategy. Finally, we introduce a soft label guided loss function to address the difficulty of backward propagation with top-$k$ selection, which could benefit future research.
Appendix A Details of QA Datasets
A.1 MMLU-Med
The Massive Multitask Language Understanding (MMLU) benchmark [15] evaluates the multitask learning capability of language models. While the full MMLU dataset encompasses 57 different tasks, we specifically extracted the medical questions for our tests, totaling 1089 questions.
A.2 MedQA-US
MedQA [53] is a multiple-choice QA dataset derived from professional medical board exams. It is available in Simplified Chinese, Traditional Chinese, and English. For our experiments, we used 1273 questions from the English version.
A.3 MedMCQA
MedMCQA [35] comprises a large number of questions from the Indian medical entrance exam, covering 2400 healthcare topics and 21 medical subjects. For the MIRAGE benchmark, we utilized the “dev” set of the original MedMCQA dataset.
A.4 PubMedQA*
PubMedQA is a research QA dataset in the biomedical field, consisting of 1000 manually annotated questions constructed from PubMed abstracts. In the MIRAGE benchmark, the reference contexts were removed. We selected a subset of 500 questions, which we refer to as PubMedQA*.
A.5 BioASQ-Y/N
BioASQ [47] is an annual biomedical QA competition. For the MIRAGE benchmark, we selected only the Machine Reading Comprehensive Track (Task B), focusing on 618 questions from recent years (2019-2023).
Appendix B Exact versions of LLMs
The exact versions of the backbone LLMs tested are listed in Table 4.
Table 4: Versions of the backbone LLMs
LLM name LLM version LLM site
ChatGPT gpt-3.5-turbo-16k https://platform.openai.com/docs/models/gpt-3.5-turbo
Llama3 Meta-Llama-3-8B https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
InternLM2 internlm2-123b https://www.sensetime.com/en/news-detail/51167237
ChatGLM3 chatglm3-6b https://huggingface.co/THUDM/chatglm3-6b
Qwen1.5 Qwen1.5-MoE-A2.7B https://huggingface.co/Qwen/Qwen1.5-MoE-A2.7B-Chat
Appendix C Experiment on soft labels
In this section, we evaluate the performance of building soft labels using different methods (TF-IDF [38], RoBERTa [32], and hitrate score). For this experiment, we fix one retriever (BM25) and the backbone LLM (Qwen1.5), then build the soft labels using these three different methods. Different routers are trained with these different soft labels and tested on various datasets of the Medical Question-Answering task. The results are grouped in Table 5.
The results indicate that there is no universal best method for building soft labels. For instance, “hitrate score” is the most effective when training on MedMCQA, while RoBERTa shows advantages on PubMedQA. Overall, the soft labels built with RoBERTa demonstrate good performance across the board. Therefore, for the remainder of the experiments, RoBERTa is adopted as the default method for building soft labels.
Table 5: The performance of the routers trained with the soft labels created with different methods, best method marked in bold.
Training datasets Methods MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
MedQA RoBERTa 0.4265$\pm$0.04 0.3711$\pm$0.04 0.3831$\pm$0.02 0.5000$\pm$0.06 0.6234$\pm$0.06 0.4608
hitrate 0.4779$\pm$0.04 0.4088$\pm$0.04 0.4330$\pm$0.02 0.4194$\pm$0.06 0.6753$\pm$0.05 0.4829
TF-IDF 0.4559$\pm$0.04 0.3019$\pm$0.04 0.4023$\pm$0.02 0.4355$\pm$0.06 0.5714$\pm$0.06 0.4334
MedMCQA RoBERTa 0.4485$\pm$0.05 0.3459$\pm$0.04 0.4023$\pm$0.02 0.4839$\pm$0.06 0.5195$\pm$0.06 0.4400
hitrate 0.4118$\pm$0.04 0.3774$\pm$0.04 0.3678$\pm$0.02 0.3387$\pm$0.06 0.5065$\pm$0.06 0.4004
TF-IDF 0.4412$\pm$0.04 0.3899$\pm$0.04 0.4119$\pm$0.02 0.5161$\pm$0.06 0.5844$\pm$0.06 0.4687
PubMedQA RoBERTa 0.4485$\pm$0.04 0.4528$\pm$0.04 0.3966$\pm$0.02 0.4677$\pm$0.06 0.6623$\pm$0.05 0.4856
hitrate 0.4118$\pm$0.04 0.4088$\pm$0.04 0.3851$\pm$0.02 0.4355$\pm$0.06 0.6104$\pm$0.06 0.4503
TF-IDF 0.4044$\pm$0.04 0.3208$\pm$0.04 0.3908$\pm$0.02 0.4677$\pm$0.06 0.5584$\pm$0.06 0.4284
Appendix D Impact of the Number of Candidate Snippets
In the previous experiment, three candidate snippets were retrieved from each granularity level of the external knowledge database. In this section, we investigate the effect of varying the number of candidate snippets on the overall performance of the RAG system. The rationale behind this exploration lies in the potential limitation of a small pool of candidate snippets, such as keeping only three. In such case, valuable snippets may be overlooked. For instance, a snippet ranked fourth or fifth in each retrieval may appear repeatedly across different granularity levels and thus might be selected after its relevance scores adjusted with weights assigned by the router. However, despite potentially high relevance, such snippets are excluded early in the retrieval process.
In this experiment, we test the RAG system equipped with MoG using different numbers of candidate snippets $k_{r}$ ($k_{r}\in$3, 8, 16, 32}). For simplicity, the backbone LLM was fixed as Qwen1.5. The experiment results are shown in Figure 5.
Generally, RAG performance improves as $k_{r}$ increases, indicating the presence of helpful knowledge in the retrieved snippets. However, too many irrelevant snippets mislead the inference of LLM based on its knowledge. MoG stands out with its high initial performance even at low $k_{r}$ values ($k_{r}$<=8 in this case), thanks to its multi-granularity filtering, which efficiently selects relevant snippets. This enhances the LLM’s accuracy without requiring a large snippet pool. MoG even strikes a balance at high $k_{r}$ values ($k_{r}$>=16 in this case) due to its threshold limiting of the different corpora, avoiding the pitfalls of excessive information retrieval. MoG’s consistently high performance demonstrates its superiority in optimizing the number of candidate snippets for effective medical question-answering.
Figure 5: Accuracy of Medical Question-Answering task with different number of candidate snippets
Appendix E Choice of Retriever
In the previous experiment, BM25 [39] was used as the retriever because it is a lightweight and popular choice in practice. In this section, we replace BM25 with other popular retrievers to evaluate their performance. Referring to the setup in MedRAG [51], we select a general-domain semantic retriever called Contriever [18], a scientific-domain retriever called SPECTER [7], and a biomedical-domain retriever called MedCPT [23]. Additionally, the Reciprocal Rank Fusion (RRF) method [9] is utilized to combine results from different retrievers, including RRF-2 (fusion of BM25 and MedCPT) and RRF-4 (fusion of all four retrievers). From Table 6, we can observe the following: Each retriever has a well-performing QA set, which results in minimal overall differences among these retrievers. Considering these conclusions, we decide to continue using BM25 for all other experiments.
Table 6: Performance of different retrievers
Corpus Retriever MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
Textbooks BM25 0.6139$\pm$0.04 0.3961$\pm$0.04 0.3907$\pm$0.02 0.3100$\pm$0.06 0.5000$\pm$0.05 0.4421
Contriever 0.5644$\pm$0.03 0.4235$\pm$0.03 0.4074$\pm$0.01 0.2500$\pm$0.04 0.5161$\pm$0.04 0.4323
SPECTER 0.5941$\pm$0.03 0.4118$\pm$0.03 0.3919$\pm$0.01 0.1900$\pm$0.04 0.4839$\pm$0.04 0.4143
MedCPT 0.5347$\pm$0.04 0.4549$\pm$0.03 0.4026$\pm$0.02 0.3100$\pm$0.05 0.5565$\pm$0.04 0.4517
RRF-2 0.5396$\pm$0.03 0.4196$\pm$0.03 0.4241$\pm$0.02 0.2800$\pm$0.04 0.5403$\pm$0.04 0.4407
RRF-4 0.5495$\pm$0.04 0.4392$\pm$0.03 0.4397$\pm$0.02 0.2600$\pm$0.04 0.6129$\pm$0.04 0.4603
Appendix F Performance of MoG on Medical QA Task (Other results)
In this section, we present the experiment result of MoG on the Medical QA task with four different corpora as the training datasets for the router. The results are grouped in Table 7.
Table 7: Accuracy of Medical Question-Answering task with MoG (trained with different corpora)
LLM Method Training Corpora (Avg Acc.)
Textbooks StatPearls PubMed Wikipedia
GLM3 CoT 0.3908 0.3908 0.3908 0.3908
MedRAG 0.4269 0.4343 0.4716 0.4582
MoG 0.4492 0.4465 0.4911 0.4241
GPT-3.5 CoT 0.6571 0.6571 0.6571 0.6571
MedRAG 0.5250 0.5229 0.6322 0.5332
MoG 0.6122 0.5921 0.6795 0.6154
InternLM CoT 0.5914 0.5914 0.5914 0.5914
MedRAG 0.4223 0.4380 0.5559 0.4506
MoG 0.5800 0.5886 0.6394 0.5810
Llama3 CoT 0.6386 0.6386 0.6386 0.6386
MedRAG 0.5127 0.5133 0.6170 0.5206
MoG 0.6221 0.6415 0.6394 0.6328
Qwen1.5 CoT 0.3845 0.3845 0.3845 0.3845
MedRAG 0.4399 0.4623 0.4499 0.4391
MoG 0.6129 0.4622 0.5145 0.4547
Appendix G Analysis on Samples Improved or Degraded
We counted the number of samples of four categories before and after applying MoG. The four categories are 1. improved (CoT gives wrong answer, MoG gives correct answer); 2. degraded (CoT gives correct answer, MoG gives wrong answer); 3. remain_correct (both CoT and MoG give correct answers); 4. remain_wrong (both CoT and MoG give wrong answers). The results are visualized as Figure 6. In this presented figure, the MoG is trained on the Wikipedia corpus. The pattern observed from this figure is similar to the ones shown in other test results. The zeroes in bars named “cot_err” and “mog_err” indicate that all the responses are correctly parsed when counting.
Around 10% of the degraded samples are randomly chosen and manually verified. We confirmed that in most (95%) degraded cases being verified, all the following statements are true:
  • several candidate snippets are retrieved correctly;
  • top-2 candidate snippets are correctly selected with the weights calculated by the router;
  • the prompt is correctly augmented and passed to backbone LLM;
  • the LLM generates the response without error;
  • the final choice (A/B/C/D or Yes/No) was correctly parsed.
In other words, the LLM changed its answer based on the snippet retrieved, it is infected by the introduced noises. The degradation is caused by the fact that the tested RAG system lacks a noise filtering mechanism, rather than by the default of MoG or MoGG.
Figure 6: Number of samples improved or degraded after application of MoG
Appendix H Performance of MoGG on Medical QA Task (Other results)
In this section, we present the experiment result of MoGG on the Medical QA task with StatPearls as the training corpus for the router. The results are grouped in Table 8.
Table 8: Accuracy of Medical Question-Answering task with MoGG (trained with StatPearls), best results marked in bold.
LLM Method MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
GLM3 CoT 0.4901$\pm$0.04 0.3294$\pm$0.03 0.3799$\pm$0.02 0.3900$\pm$0.05 0.5645$\pm$0.04 0.4181
MedRAG 0.4901$\pm$0.04 0.3294$\pm$0.03 0.3799$\pm$0.02 0.4400$\pm$0.05 0.5323$\pm$0.04 0.4343
MoG 0.4851$\pm$0.04 0.2902$\pm$0.03 0.3799$\pm$0.02 0.4400$\pm$0.05 0.6371$\pm$0.04 0.4465
MoGG 0.4802$\pm$0.04 0.3059$\pm$0.03 0.3668$\pm$0.02 0.4600$\pm$0.05 0.5806$\pm$0.04 0.4387
GPT-3.5 CoT 0.7624$\pm$0.03 0.6706$\pm$0.03 0.5866$\pm$0.02 0.4200$\pm$0.05 0.7339$\pm$0.04 0.6347
MedRAG 0.6881$\pm$0.03 0.6549$\pm$0.03 0.4552$\pm$0.02 0.2600$\pm$0.04 0.5565$\pm$0.04 0.5229
MoG 0.7277$\pm$0.03 0.6471$\pm$0.03 0.5400$\pm$0.02 0.3200$\pm$0.05 0.7258$\pm$0.04 0.5921
MoGG 0.7624$\pm$0.03 0.6314$\pm$0.03 0.5460$\pm$0.02 0.3700$\pm$0.05 0.7177$\pm$0.04 0.6055
InternLM CoT 0.7228$\pm$0.03 0.6000$\pm$0.03 0.5352$\pm$0.02 0.3700$\pm$0.05 0.7339$\pm$0.04 0.5924
MedRAG 0.6584$\pm$0.03 0.5137$\pm$0.03 0.3596$\pm$0.02 0.1500$\pm$0.04 0.5081$\pm$0.04 0.4380
MoG 0.6881$\pm$0.03 0.5961$\pm$0.03 0.5448$\pm$0.02 0.3800$\pm$0.05 0.7339$\pm$0.04 0.5886
MoGG 0.6782$\pm$0.03 0.5725$\pm$0.03 0.3500$\pm$0.02 0.7500$\pm$0.05 0.7500$\pm$0.04 0.6201
Llama3 CoT 0.7277$\pm$0.03 0.6392$\pm$0.03 0.5663$\pm$0.02 0.5900$\pm$0.05 0.7177$\pm$0.04 0.6482
MedRAG 0.6089$\pm$0.03 0.5882$\pm$0.03 0.4170$\pm$0.02 0.3800$\pm$0.05 0.5242$\pm$0.04 0.5133
MoG 0.7376$\pm$0.03 0.5961$\pm$0.03 0.5317$\pm$0.02 0.6000$\pm$0.05 0.7419$\pm$0.04 0.6415
MoGG 0.7030$\pm$0.03 0.6275$\pm$0.03 0.5436$\pm$0.02 0.5900$\pm$0.05 0.7258$\pm$0.04 0.6380
Qwen1.5 CoT 0.4604$\pm$0.04 0.3255$\pm$0.03 0.3883$\pm$0.02 0.2000$\pm$0.04 0.5484$\pm$0.04 0.3845
MedRAG 0.5495$\pm$0.04 0.4471$\pm$0.03 0.4146$\pm$0.02 0.3600$\pm$0.05 0.5403$\pm$0.04 0.4623
MoG 0.5446$\pm$0.04 0.3882$\pm$0.03 0.4337$\pm$0.02 0.3400$\pm$0.05 0.6048$\pm$0.04 0.4623
MoGG 0.5446$\pm$0.04 0.4157$\pm$0.03 0.4253$\pm$0.02 0.3500$\pm$0.05 0.6290$\pm$0.04 0.4729