### v2

# Mix-of-Granularity: Optimize the Chunking Granularity for Retrieval-Augmented Generation

Abstract
Integrating information from various reference databases is a major challenge for Retrieval-Augmented Generation (RAG) systems because each knowledge source adopts a unique data structure and follows different conventions. Retrieving from multiple knowledge sources with one fixed strategy usually leads to under-exploitation of information. To mitigate this drawback, inspired by Mix-of-Expert, we introduce Mix-of-Granularity (MoG), a method that dynamically determines the optimal granularity of a knowledge source based on input queries using a router. The router is efficiently trained with a newly proposed loss function employing soft labels. We further extend MoG to MoG-Graph (MoGG), where reference documents are pre-processed as graphs, enabling the retrieval of distantly situated snippets. Experiments demonstrate that MoG and MoGG effectively predict optimal granularity levels, significantly enhancing the performance of the RAG system in downstream tasks. The code of both MoG and MoGG are released in https://github.com/ZGChung/Mix-of-Granularity.
Mix-of-Granularity: Optimize the Chunking Granularity for Retrieval-Augmented Generation
Zijie Zhong Shanghai AI Laboratory zhongzijie@pjlab.org.cn                        Hanwen Liu Beihang University liuhwen@buaa.edu.cn                        Xiaoya Cui Beihang University xiaoya@buaa.edu.cn
Xiaofan Zhang [^1] Corresponding authors Shanghai AI Laboratory zhangxiaofan@pjlab.org.cn                        Zengchang Qin [^1] Corresponding authors Beihang University, theSight Technology zcqin@buaa.edu.cn
1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) has become a popular method for enhancing Large Language Models (LLMs). The core concept of RAG involves retrieving relevant information from external knowledge bases to provide additional context to the LLM, enabling it to generate more precise and grounded responses. RAG offers a promising and practical solution to mitigate LLMs’ hallucinations because (1) it can be applied to any LLM, even those accessible only via APIs, and (2) the reference information is easy to modify or update. Many LLM-based products are supported by RAG systems, with examples spanning various industries such as customer service, advertising and marketing, education and e-learning, healthcare, and e-commerce and retailing Khan (2023); Xiong et al. (2024); Soman and Roychowdhury (2024); Ke et al. (2024); Radeva et al. (2024).
The quality of the retrieved snippets is crucial for the final generation, consequently, much research has focused on the retrieval phase. Currently, most RAG systems follow the Dual-Encoder Architecture Dong et al. (2022) (DEA) paradigm, in which the reference documents are divided into small snippets (chunks), encoded by specific encoders, and then stored in the vector database (e.g. FAISS Johnson et al. (2017) or Neo4j Company (2012)) as embeddings. Thanks to its scalability, the DEA paradigm shows great potential for connecting LLMs with knowledge database of different formats, including knowledge graphs, textbooks, or Wikipedia articles.
Figure 1: MoG automatically selects the optimal granularity when extracting information from the reference database (scenarios 1 and 2), achieving both high pertinence and coverage. When relevant information is dispersed across distant sections (scenario 3), the reference documents are pre-processed as graph, then MoGG is applied to retrieve these separate snippets from the best hopping range.
Optimizing the chunk size of the reference knowledge database is essential for enhancing the precision and recall during the retrieval phase. However, this optimization presents significant challenges in the implementation of RAG systems for two primary reasons: (1) An optimal chunking size needs to be determined for each knowledge database due to their dissimilar data structures and information densities. For example, the optimal chunking size for medical textbooks is longer than the one for Medical Knowledge Graph Database (MKGD like Hetionet Himmelstein et al. (2017)), as Medical textbooks typically contain lengthy passages, while an MKGD consists of entities represented by shorter terms. (2) Even within a single knowledge database, using a uniform chunking size for all input queries can yield suboptimal retrieval results, as the queries themselves exhibit varying levels of granularity. As shown in Figure 1), when the user asks about one disease (fine-grained question), chunking the reference document in finer granularity is better; whereas, when the user asks for broader information (coarse-grained question), a more coarse granularity is preferred. In practice, the optimal "uniform chunking size" is determined through parameter tuning, which is not only tedious but also does not guarantee high precision and recall.
Therefore, the community calls for a method to dynamically determine the optimal chunking size, for which we propose the Mix-of-Granularity (MoG). We draw inspiration from Mix-of-Experts Chen et al. (2022), which is a machine learning architecture that dynamically selects the most pertinent “expert sub-network” for each input token using a router module. Similarly, MoG involves leveraging a router to select the best granularity levels from which the reference snippets are extracted.
Even with this flexibility newly introduced, MoG still has difficulty dealing with broader queries that require distantly situated snippets, e.g. snippets stored in different knowledge databases. In such cases (cross-document question in Figure 1), adjusting only the granularity is not helpful because the necessary information is so distantly located that it can never be covered by a reasonably large chunking window. To better answer these broader queries, we extend MoG to MoG-Graph (MoGG). In MoGG, the reference documents are pre-processed as a graph, allowing relevant snippets to be included as neighbors of each other, regardless of their distance in the original databases. This extension further improves the performance of cross-source retrieval.
When training MoG(G) under the supervised learning setting, the backward propagation is blocked by the top-$k$ selection at the end of the retrieval phase, which is a common practice in most RAG systems. To solve it, we introduce a loss function using soft labels. Soft labels are approximate training signals generated using offline algorithms or models like TF-IDF Ramos (2003) or RoBERTa Liu et al. (2019). With the soft labels, the top-$k$ selection is excluded from the training process, thus the issue of backward propagation is circumvented and the training is accelerated.
In conclusion, the main contributions of this work are:
(1) We propose MoG, which dynamically determines the optimal granularity level for retrieval with the help of a router, achieving a balanced trade-off between precision and recall.
(2) We extend MoG to MoGG by reorganizing the reference document in the form of a graph, thereby further improving the quality of cross-source retrieval involving multiple knowledge databases.
(3) We introduce a loss function utilizing soft labels to overcome the challenges associated with training with top-$k$ selection.
2 Related Work
2.1 Retrieval-Augmented Generation
RAG Lewis et al. (2020) has emerged as a standard practice to enhance the LLMs by mitigating their problems of “hallucinations” and knowledge cut-off. A RAG system typically includes a Retriever that extracts relevant information from an external knowledge database, and a backbone LLM to generate grounded responses by in-context learning Dong et al. (2023). Previous retrieval-focused methods have evolved from retrieving tokens Khandelwal et al. (2020) or entities Nishikawa et al. (2022) to more complex structures like chunks Ram et al. (2023) or graphs Kang et al. (2023). Granularity matters a lot in retrieval, as coarse-granularity-retrieval yields more information with lower precision, while fine-granularity-retrieval offers comprehensive information at the cost of efficiency. More strategies like single Wang et al. (2023b); Shi et al. (2023), adaptive Jiang et al. (2023); Huang et al. (2024), or multiple retrieval Izacard et al. (2022b) are introduced to improve the retrieval phase’s performance. Regarding the generation phase, various information fusion techniques are developed to integrate retrieval results to LLM in its input Khattab et al. (2023), intermediate Borgeaud et al. (2022), or output layers Liang et al. (2024).
2.2 Chunking Optimization
Optimal chunk size is crucial for RAG system as breaking down documents into small chunks is the first step to encode them. Naive chunking strategies, such as “fixed-size” chunking and “recursive chunking,” attempt to create snippets of identical size. Later works explored more chunking optimization techniques. One line of work focuses on increasing the recall of retrieval. For example, the “Sliding Window Chunking” Safjan (2023) allows layered retrieval by merging globally related information across multiple processes. The “Parent Document Retrieval” team (2023) retrieves using small chunks and returns larger blocks of context for later generation. Another line of work seeks to include more semantic information of the context to improve retrieval accuracy. “Metadata Filtering” Siegler (2024) leverages document metadata to filter snippets; “Context-Enriched Chunking” team (2024) breaks down information into segments and adds semantic summaries before retrieval; while “Windowed Summarization Chunking” team (2024) enriches each chunk with a windowed summary of the previous chunks. Many of these techniques, including MoG(G), are compatible with each other and can be combined to achieve better performance.
2.3 Graph-Based Text Processing
Graph-based text processing techniques combine research in graphs and text retrieval. Previous works exploit semantic similarities between small snippets (a sentence or several words) and reorganize the text material into a graph using Entity Recognition and Relation Construction algorithms Melnyk et al. (2022); Guo et al. (2020); You et al. (2018). Breaking the constraint of the single dimension of text corpus, these methods allow chunks of the same topic to be grouped as neighbors in a graph, thus show great potential in tasks requiring long context reasoning or multi-hop reasoning. For example, “graph indexing” Gao et al. (2024) transforms entities and relationships into nodes and connections, improving the relevance of retrieved snippets significantly. RAPTOR Sarthi et al. (2024) organizes snippets as a tree (a special form of a graph) by recursively clustering them, where all non-leaf nodes correspond to summaries of their child nodes. This processing allows access to information at different granularity levels, which is beneficial to summarization tasks. In GMoE Wang et al. (2023a), authors use different expert networks to handle hop-$1$, hop-$2$, and mixed hop-$1$ hop-$2$ neighbors of a node in a graph, which inspired our design of MoGG.
3 Methodology
3.1 Preliminaries
MoG(G) is designed to enhance the Retriever of a RAG system. A typical RAG system comprises a Retriever $\mathcal{R}$, a Generator $\mathcal{G}$, and a series of external reference corpus $\mathcal{K}=\{\mathcal{K}_{1},\mathcal{K}_{2},...,\mathcal{K}_{k}\}$. A RAG system takes a user query $q$ as input, retrieves reference document pieces (snippets or chunks) from $\mathcal{K}$, and uses these snippets to help the $\mathcal{G}$ produce the final responses. A popular architecture for the Retriever is DEA Dong et al. (2022), where the query $q$ and all the snippets in $\mathcal{K}$ are encoded into embeddings ($e_{q}$ and $e_{s}$) using the same encoder $\mathcal{E}$. The extraction of relevant snippets is achieved by calculating the similarity between $e_{q}$ and $e_{s}$. The snippets with highest similarity scores are extracted and injected into the backbone LLM via prompt.
3.2 Naive MoG
3.2.1 Multi-granularity Router
We apply the idea of Mix-of-Expert Chen et al. (2022) (MoE) to automatically determine the best granularity level in the retrieval phase. In a MoE system, different input tokens are routed to the best expert network based on the weights output by the router. Similarly, in MoG, a router is trained to predict the importance weight of different granularity levels based on the user’s input, so that the snippets from the best granularity level are prioritized. By employing such a routing optimization method, we can effectively adjust the chunk size according to different scenarios. Take one corpus $\mathcal{K}$ as example, its documents are chunked in $n_{gra}$ candidate granularities before the retrieval (in Figure 2, $n_{gra}$ = 5). Although some methods can dynamically adjust chunks using well-trained models, we prioritize the most commonly used and simplest chunking method from the perspective of practical efficiency. The snippets are chunked without overlap, and each chunk in $j$ ($j\in[2,n_{gra}]$) granularity level is formed with 2 adjacent chunks in $j-1$ granularity level (granularity level 1 is the finest). Each chunk is assigned a similarity score using BM25 Robertson and Zaragoza (2009) with respect to user’s input query $q$. In each granularity, $k_{r}$ most relevant snippets are extracted ($k_{r}$ = 3 in Figure 2), forming a pool of candidate snippets of size $n_{gra}\times k_{r}$. In parallel, $q$ is encoded via RoBERTa before being mapped to a weight $w$ (a vector of the same length as $n_{gra}$) by the router. The chunks’ similarity scores are then weighted and summed with $w$.
Empirically, directly retrieving the top snippets across different granularities based on their weighted similarity scores will give biased results, because the scores of more coarsed granularity levels are systematically higher (more discussions in Appendix A). Therefore, we adopt another selection strategy involving two steps: (1) select the top-$k$ ($k$ = 1 in Figure 2) finest-grained chunk $chunk_{r}$; (2) retrieve the chunk containing $chunk_{r}$ from the optimal granularity level $g_{r}$. The selection process can be formalized as follows:
$chunk_{r}$ $=\underset{c\in\mathcal{C}}{\operatorname{top_{k}-argmax}}(t_{rs}(c)\cdot w);$ (1)
$g_{r}$ $=\underset{g\in[1,n_{gra}]|w_{g}\neq 0}{\operatorname{argmax}}w_{g}\;$ (2)
where $\mathcal{C}$ represents the set of all chunks of the finest granularity level, $t_{rs}(.)$ represents the relevance score of a chunk, and $w_{g}$ represents the $g$-th element in $w$.
Two snippets are highlighted in Figure 2 as examples to help understanding. For each chunk of the finest granularity level, if it is retrieved among the top-$k_{r}$ snippets from one granularity level, its relevance score for that level is recorded as the actual relevance score (like the red chunk); otherwise, it is filled with 0 (like some zeroes padded for the blue chunk). In this way, we get the relevance scores $t_{rs}$ of each finest snippet. The blue chunk has highest weighted similarity score, it is therefore selected as $chunk_{r}$. In this example, the final snippet selected is the one extracted from 3rd granularity level containing $chunk_{r}$.
Figure 2: MoG mechanism prioritizes the chunks retrieved from optimal granularity level, which is determined by the router based on the user input query.
3.2.2 Soft-labels
In MoG, we train a Multi-Layer Perceptron (MLP) as the router in a supervised learning method. The input for the router is the embedding of the input query $q$ generated with RoBERTa Liu et al. (2019), which is then mapped to $w$ by the router. The intrinsic training signal is the “labels” ($l$) in the Medical Question-Answering (MQA) datasets. In some MQA datasets, the ground truth snippets to be retrieved are provided, so we use them as “labels” directly; otherwise, the concatenation of the strings of the “Question” and “Answer” is used as the “label”. A natural training objective is to maximize the semantic similarity between $chunk_{r}$ and $l$ by adjusting $w$. Such a label helps us to train the router to choose the best granularity based on the input query, which is a prior, with solid posterior information (the similarity between the ans and the text of different granularities). Unfortunately, this label can not be used to guide the training directly because there is a non-differentiable top-$k$ selection in the way. The soft labels are proposed to bypass the top-$k$ selection during the training: For each query $q$, the most relevant snippet ($S_{best}$) is retrieved from the reference documents of each granularity level with BM25 Robertson and Zaragoza (2009). The semantic similarity between each snippet in $S_{best}$ and the label $l$ is then calculated (with static models including TF-IDF Ramos (2003), RoBERTa Liu et al. (2019), or hitrate score) and stored in $sim_{best}$. We create a soft label of 0.8 (resp. 0.2) for the most (resp. the second) similar snippet in $S_{best}$, and pad 0 for the other snippets.
For example, the soft labels corresponding to $sim_{best,1}$ [0, 0.32, 0.11, 0.88, 0.45] and $sim_{best,2}$ [0.95, 0.07, 0.22, 0.11, 0.19] are $sl_{1}$ [0, 0, 0, 0.8, 0.2] and $sl_{2}$ [0.8, 0, 0.2, 0, 0], respectively.
The values of the soft labels are designed to guide the router in distinguishing the relative importance among the granularity candidates. Empirically, setting these values to either [0.8, 0.2, 0] or [0.7, 0.3, 0] yields similar results. With the soft labels ($sl$) built, we can train the router to predict a high value (0.8) for the optimal granularity level, while conserving certain flexibility to choose the second-best granularity level. The router is trained by minimizing a Binary Cross Entropy loss function ($l_{bce}$):
$l_{bce}=\sum_{i\in len(w)}-[sl_{i}\cdot\log(w_{i})+(1-sl_{i})\cdot\log(1-w_{i})].$ (3)
3.3 MoGG: MoG with Graph-context
With MoG, the adjacent snippets with relevant knowledge can be retrieved altogether by adjusting the granularity level. This method is particularly effective when information centered around the same topic is stored in adjoining sentences. However, in most cases, answering a complex question requires reasoning over information stored in different paragraphs or even different documents. A common solution is to perform more retrievals at a finer granularity level, retrieving only highly relevant small pieces to form a comprehensive reasoning chain. This approach is inconvenient because determining the optimal number of retrievals often involves manually adjusting $k$, which is challenging because the whole tuning process is time-consuming and $k$ can not go infinitely large.
To overcome this challenge, an intuitive approach is to reorganize the reference document and group the relevant information together. Motivated by this, we propose a more applicable framework, MoGG, by enhancing MoG with a preprocessing step that organizes the documents in $\mathcal{K}$ as a graph. As illustrated in Figure 3, each document is initially split into small pieces consisting of one or two sentences, and each piece is treated as a separate node in the graph. To determine the edges in the graph, an index is first created with all these nodes. Then, each node is used as a “query” to search for the $k_{graph}$ (set as 3 in Figure 3) most relevant nodes using BM25. An edge is then added between two nodes if the similarity between them meets a predefined threshold $T_{graph}$.
Figure 3: Pre-processing the reference document to form graphs. The concept of “granularity level” is changed into “hopping range” in graphs.
This pre-processing extends the concept of “context” in linear text to “neighbors” in a graph. Also, granularity levels are adapted to “hopping ranges”: in a non-graph setting, a larger granularity level corresponds to more adjacent sentences being grouped in a chunk; in a graph setting, a larger granularity level corresponds to nodes within a larger hopping range of a centered node being grouped as one chunk. To avoid context redundancy, duplicate nodes in the neighbors are considered only once. Similar to MoG, the documents in the external knowledge database $\mathcal{K}$ are chunked with $n_{gra}$ different hopping ranges and then encoded into embeddings, while the rest of the MoG remains unchanged. By transforming the documents into a graph and defining granularity based on hopping ranges, MoGG effectively captures dispersed relevant information, allowing for more comprehensive and efficient retrieval.
4 Experiments
4.1 Corpus and Medical QA datasets
A reference knowledge database used in a RAG system is often termed a “corpus”. To form the corpora, data from various sources were collected, including the widely-used PubMed Galileo Mark Namata and Huang (2012) corpus for all biomedical abstracts, the StatPearls Publishing (2024) corpus for clinical decision support, the TextbooksJin et al. (2020) corpus covering medical textbook knowledge, and the Wikipedia Wikimedia Foundation (2024) corpus for general knowledge. The four corpora are combined to form a larger corpus, named “MedCorp”. For each corpus used in our experiments, there are 5 granularity levels tested. The chunking size of the second granularity level is set to be the same as the one in MedRAG to facilitate the comparison of the results. The chunking size of the {1st, 3rd, 4th, 5th} granularity level is set to be {1/2, 2, 4, 8} times of the one in MedRAG.
The performance of the RAG system are evaluated using five MQA datasets following MIRAGE benchmark Xiong et al. (2024): MMLU-Med Hendrycks et al. (2021), MedQA-US Zhang et al. (2018), MedMCQA Pal et al. (2022), PubMedQA* Jin et al. (2019), and BioASQ-Y/N Tsatsaronis et al. (2015). Specifically, only the biomedical questions are kept and the ground truth supporting contexts are removed during testing.
The detailed descriptions of each copora and each QA dataset, as well as their important statistics, are included in Appendix B. We have begun to test our method using medical question-answering datasets, as they represent a knowledge-intensive domain. We posit that significant improvements demonstrated by the tests on this knowledge-intensive field suggest MoG’s potential effectiveness in other domains with lower knowledge dependencies and higher error tolerances.
4.2 Experiment Setup
All backbone LLMs, whether accessed via API or local deployment, are run under off-the-shelf settings. The exact versions of the backbone LLMs are listed in Appendix C. Experiments are conducted on Nvidia GeForce 3090 and 4090 GPUs. The code is written using the PyTorch framework, utilizing an Adam optimizer with a learning rate of 0.001. Each training job is run until the convergence of the loss value. During the experiment, there are two top-k selections. Unless specified, when retrieving snippets from each corpus, we select the top-3 snippets; when all the snippets are retrieved from each corpus, we select the top-2 snippets with the highest relevance scores to pass to the backbone LLM. When snippets are too long, they are truncated automatically to fit the LLM’s context window size. The router requires approximately 12GB of GPU memory for training and 6GB for inference. Utilizing a caching mechanism, we efficiently completed 35 training sessions and over a hundred inferences, each training session taking around 4 hours for 1000 epochs.
4.3 Performance of MoG on MQA Task
As mentioned above, we test the effectiveness of MoG on MQA datasets. For each question, the RAG system is tasked with choosing the best answer(s) from the given options. The performance of the entire RAG system is measured by the Exact Matching accuracy of the answers. To prevent knowledge leakage, only the question is used (options excluded) to retrieve reference documents from the external knowledge database. The router of MoG guides the retrieval system to choose the optimal granularity. Qualitatively speaking, when tested on different datasets, the router trained with Textbook corpus shows a preference for different granularity levels. For instance, on the PubMedQA dataset, the finest granularity snippets are selected most frequently. This is because the questions in PubMedQA are typically precise and can be answered with short reference snippets. An example result figure is shown in Figure 4, with a more detailed discussion included in Appendix D.
Figure 4: Averaged weights of different granularity levels on different QA datasets
LLM Method MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
GLM3 CoT 0.4356$\pm$0.04 0.3451$\pm$0.03 0.3250$\pm$0.02 0.3400$\pm$0.05 0.5081$\pm$0.04 0.3908
MedRAG 0.4950$\pm$0.04 0.3804$\pm$0.03 0.3632$\pm$0.02 0.5100$\pm$0.05 0.6532$\pm$0.04 0.4804
MoG 0.5198$\pm$0.04 0.3529$\pm$0.03 0.3716$\pm$0.02 0.5400$\pm$0.05 0.6774$\pm$0.04 0.4923
GPT-3.5 CoT 0.7525$\pm$0.03 0.6118$\pm$0.03 0.5890$\pm$0.02 0.5200$\pm$0.05 0.7339$\pm$0.04 0.6571
MedRAG 0.6931$\pm$0.03 0.6510$\pm$0.03 0.4671$\pm$0.02 0.6000$\pm$0.05 0.8306$\pm$0.03 0.6484
MoG 0.7129$\pm$0.03 0.6471$\pm$0.03 0.5532$\pm$0.02 0.6200$\pm$0.05 0.7823$\pm$0.04 0.6631
InternLM CoT 0.7426$\pm$0.03 0.6118$\pm$0.03 0.5269$\pm$0.02 0.3500$\pm$0.05 0.7258$\pm$0.04 0.5914
MedRAG 0.6040$\pm$0.04 0.5294$\pm$0.03 0.3847$\pm$0.03 0.4300$\pm$0.05 0.7661$\pm$0.04 0.5428
MoG 0.7129$\pm$0.03 0.5961$\pm$0.03 0.5436$\pm$0.02 0.4100$\pm$0.05 0.7661$\pm$0.04 0.6057
Llama3 CoT 0.7079$\pm$0.03 0.6431$\pm$0.03 0.5663$\pm$0.02 0.5500$\pm$0.05 0.7258$\pm$0.04 0.6386
MedRAG 0.6040$\pm$0.03 0.5725$\pm$0.03 0.4313$\pm$0.02 0.5600$\pm$0.05 0.7823$\pm$0.04 0.5900
MoG 0.7228$\pm$0.03 0.6000$\pm$0.03 0.5627$\pm$0.02 0.6400$\pm$0.05 0.7984$\pm$0.04 0.6648
Qwen1.5 CoT 0.4604$\pm$0.04 0.3255$\pm$0.03 0.3883$\pm$0.02 0.2000$\pm$0.03 0.5484$\pm$0.04 0.3845
MedRAG 0.5594$\pm$0.03 0.4353$\pm$0.03 0.4038$\pm$0.02 0.3400$\pm$0.05 0.5403$\pm$0.04 0.4558
MoG 0.5941$\pm$0.04 0.4235$\pm$0.03 0.4301$\pm$0.02 0.4700$\pm$0.05 0.6694$\pm$0.04 0.5174
Table 1: Accuracy of Medical Question-Answering task with MoG (trained with MedCorp), best results in bold.
MoG is integrated into one same RAG system, with which we conduct the MQA task. Backbone LLMs are altered to cover some of the popular ones, such as ChatGPT Brown et al. (2020), InternLM2 Cai et al. (2024), Llama3 AI (2024), GLM3 Du et al. (2022), and Qwen1.5 Bai et al. (2023). The router is trained with the soft labels built using RoBERTa Liu et al. (2019), this choice is justified by the experiment in Appendix E. To investigate the effect of varying the number of candidate snippets on RAG system performance, we conducted experiments with different snippet counts (details are provided in Appendix F). The retriever is fixed as BM25 Robertson and Zaragoza (2009), with a further discussion on the performance of different retrievers included in Appendix G. In Table 1 we present the results obtained with the router trained with MedCorp corpus (the results obtained with routers trained on four single corpora are presented in Appendix H).The results are compared with two baselines: CoT and MedRAG. CoT baseline adopts Chain-of-Thought Wei et al. (2022) prompting and does not leverage any external knowledge database to enhance its response. MedRAG baseline is a simple RAG system with only 1 candidate granularity level introduced in MedRAG paper Xiong et al. (2024).
The results demonstrate that MoG consistently enhanced the performance of the RAG system across different backbone models when compared with MedRAG, though not necessarily better than CoT. The reason is that the RAG system we used has no noise filters or any quality control mechanism, thus the noise is injected along with the knowledge via prompts. A detailed analysis of the number of samples improved or degraded by the application of MoG is included in the Appendix I, in which we manually verified that the majority of degradation is caused by noise. We also find that MoG improves the accuracy score more when applied on weaker LLMs (like ChatGLM and Qwen), probably because they have less knowledge stored in their internal parameters and, thus could benefit more from the retrieved snippets.
LLM Method MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
GLM3 CoT 0.4356$\pm$0.04 0.3451$\pm$0.03 0.3250$\pm$0.02 0.3400$\pm$0.05 0.5081$\pm$0.04 0.3908
MedRAG 0.4802$\pm$0.04 0.3569$\pm$0.03 0.3811$\pm$0.02 0.3600$\pm$0.05 0.5565$\pm$0.04 0.4269
MoG 0.5545$\pm$0.04 0.2941$\pm$0.03 0.3548$\pm$0.02 0.4700$\pm$0.05 0.5726$\pm$0.04 0.4492
MoGG 0.5347$\pm$0.04 0.3176$\pm$0.03 0.3608$\pm$0.02 0.4500$\pm$0.05 0.5645$\pm$0.04 0.4455
GPT-3.5 CoT 0.7525$\pm$0.03 0.6118$\pm$0.03 0.5890$\pm$0.02 0.5200$\pm$0.05 0.7339$\pm$0.04 0.6571
MedRAG 0.7277$\pm$0.03 0.6745$\pm$0.03 0.4468$\pm$0.02 0.2600$\pm$0.04 0.5161$\pm$0.04 0.5250
MoG 0.7525$\pm$0.03 0.6667$\pm$0.03 0.5603$\pm$0.02 0.5200$\pm$0.05 0.7016$\pm$0.04 0.6122
MoGG 0.7673$\pm$0.03 0.6784$\pm$0.03 0.5233$\pm$0.02 0.4700$\pm$0.05 0.6452$\pm$0.04 0.6168
InternLM CoT 0.7426$\pm$0.03 0.6118$\pm$0.03 0.5269$\pm$0.02 0.3500$\pm$0.05 0.7258$\pm$0.04 0.5914
MedRAG 0.6188$\pm$0.03 0.5569$\pm$0.03 0.3118$\pm$0.02 0.1400$\pm$0.03 0.4839$\pm$0.04 0.4223
MoG 0.7277$\pm$0.03 0.5725$\pm$0.03 0.5281$\pm$0.02 0.3400$\pm$0.05 0.7339$\pm$0.04 0.5804
MoGG 0.7228$\pm$0.03 0.5882$\pm$0.03 0.5173$\pm$0.02 0.3400$\pm$0.05 0.7339$\pm$0.04 0.5804
Llama3 CoT 0.7079$\pm$0.03 0.6431$\pm$0.03 0.5663$\pm$0.02 0.5500$\pm$0.05 0.7258$\pm$0.04 0.6386
MedRAG 0.6485$\pm$0.03 0.5961$\pm$0.03 0.4146$\pm$0.02 0.3800$\pm$0.05 0.5242$\pm$0.04 0.5127
MoG 0.7228$\pm$0.03 0.6196$\pm$0.03 0.5484$\pm$0.02 0.5100$\pm$0.05 0.7097$\pm$0.04 0.6221
MoGG 0.7030$\pm$0.03 0.5961$\pm$0.03 0.5460$\pm$0.02 0.5200$\pm$0.05 0.7661$\pm$0.04 0.6262
Qwen1.5 CoT 0.4604$\pm$0.04 0.3255$\pm$0.03 0.3883$\pm$0.02 0.2000$\pm$0.03 0.5484$\pm$0.04 0.3845
MedRAG 0.5941$\pm$0.03 0.4000$\pm$0.03 0.3835$\pm$0.02 0.3300$\pm$0.05 0.4919$\pm$0.04 0.4399
MoG 0.5792$\pm$0.03 0.3843$\pm$0.03 0.4110$\pm$0.02 0.3300$\pm$0.05 0.6129$\pm$0.04 0.4635
MoGG 0.5594$\pm$0.03 0.4314$\pm$0.03 0.4480$\pm$0.02 0.3000$\pm$0.05 0.6371$\pm$0.04 0.4752
Table 2: Accuracy of Medical Question-Answering task with MoGG (trained with Textbooks), best results in bold.
4.4 Performance of MoGG on MQA Task
Similarly to the experiment of MoG, we test the performance of MoGG on MQA datasets. We tested only the performance of MoGG with routers trained on Textbooks and StatPearls corpora because training with the other two much larger corpora is too time-consuming. In Table 2 we present the results obtained when trained with Textbook corpus (results obtained with StatPearls showed in Appendix J with similar patterns). From the table, we can tell that MoGG can further improve the averaged accuracy scores. By comparing Table 1 and Table 2, we can find that, even when trained with significantly fewer samples (Textbooks corpus is a tiny subset of MedCorp corpus, accounting for only about 0.2% of all the snippets in MedCorp), MoGG brings more significant improvement in terms of the averaged accuracy score with respect to MedRAG than MoG. This finding highlights that MoGG is more efficient than MoG thanks to its flexible way of organizing the reference snippets (in the form of a graph). We conducted an ablation test with Llama 3 and the results are included in Appendix K.
There is a general performance drop in the metrics in Table 2 compared to Tabel 1 because the Textbooks corpus is only a tiny subset of the MedCorp used in Table 1). To facilitate the comparison, we conducted the same experiment in Table 1 with only the Textbooks corpus, and the detailed results are reported in Appendix L.
4.5 Execution time and storage efficiency
MoG(G) is proposed to increase the precision and recall of the retrieval phase at a reasonable cost of computational efficiency because the quality of the retrieved chunks is prioritized for application scenarios like the medical environment. We measured the average inference time with the different number of candidate granularity levels, and the results show that increasing the number of granularity levels will only increase a marginal increase in execution time. (Details in Appendix M)
In terms of storage, while additional space is required to store the embeddings of the corpus at different granularities, we only need to store one copy of the corpus and five sets of embeddings. This engineering optimization results in a space requirement of only 2.7 times the size of the original corpus. We believe this represents an acceptable overhead. Furthermore, in our RAG system specialized in the medical domain, a corpus containing about 10GB of plain text can already cover a wide range of questions. Thus, we believe this will not be an obstacle to MoG(G)’s wider application.
We believe this overhead of computing resources added is worthwhile because in most cases, MoG performs much better than MedRAG and CoT baselines. On smaller models, MoG shows an average improvement of 5% compared to MedRAG and 8.7% compared to CoT. Given the dataset size of approximately 7000, these improvements are statistically and practically significant.
5 Limitations and Broader Impacts
Our work serves as an early trial in dynamic chunking strategy, with the following major directions for improvement. (1) MoG(G)’s candidate granularity levels are manually assigned. It could be more efficient if an algorithm automatically set these granularity levels to avoid excessive grid-searching for parameter optimization. (2) Current router uses only the semantic information of the input query to predict the best granularity level. (3) Previous studies have demonstrated that the use of length normalization is crucial in information retrieval-related fields. This paper primarily focuses on applying the concept of Mix-of-Experts to this application scenario. In future work, we will also take related issues into consideration. Incorporating more information (like query type or expected response length) into the router can potentially improve the results. However, the router also introduces a new security risk: a compromised router could redirect knowledge retrieval to malicious sources, injecting incorrect or even harmful information into the backbone LLM. Therefore, it is crucial to protect and monitor the router to mitigate this risk.
6 Conclusion
In this work, we present MoG, a mechanism to dynamically choose the best granularity when retrieving information from an external knowledge database. When applied to a RAG system, MoG helps retrieve more relevant information while reducing noise. MoG is further extended as MoGG, where reference documents are pre-processed as graphs. This extension allows distantly situated information to be retrieved simultaneously, overcoming the limitations of a fixed top-$k$ selection strategy. Finally, we introduce a soft label guided loss function to address the difficulty of backward propagation with top-$k$ selection, which could benefit future research.
Acknowledgements
We would like to express our gratitude to all the supervisors and colleagues at Shanghai Artificial Intelligence Laboratory for their invaluable insights, feedback, and support throughout the research process. We also thank the reviewers for their constructive comments, which greatly improved the quality of this paper.
Appendix A More Discussion About Selection Process
The intuitive selection process would be to select the top snippets with the highest weighted similarity scores. However, this method is not adopted for the following reasons:
(1) It is biased, because more-coarsed snippets tend to have higher similarity scores, even though they might include more noise.
(2) It is not rebust to rely on the router’s output weight to control such imbalance between granularity levels, because the router is not traine for this objective. The router is trained with soft labels, which are approximate training signals that can only teach the model to distinguish the 1st and 2nd optimal granularity level from the rest.
(3) In practice, when $\mathcal{K}$ is too large, it has to be stored in separated vector databases. A common solution is use one vectorbase to store one granularity. This solution will make the similarity scores across granularity levels NOT directly comparable with each other.
Based on these reasons, we designed the selection strategy presented in the paper. The rationale behind the proposed strategy is that: Since we aim to extract the globally optimal snippet, we need a way to calculate a “global highest similarity score”. Given that the finest-grained snippets are the only ones that appear across all granularity levels following our initial chunking approach (fixed-size, non-overlapping), we calculate the weighted sum based on these finest-grained snippets.
Appendix B Details of QA Datasets
B.1 MMLU-Med
The Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) evaluates the multitask learning capability of language models. While the full MMLU dataset encompasses 57 different tasks, we specifically extracted the medical questions for our tests, totaling 1089 questions.
B.2 MedQA-US
MedQA Zhang et al. (2018) is a multiple-choice QA dataset derived from professional medical board exams. It is available in Simplified Chinese, Traditional Chinese, and English. For our experiments, we used 1273 questions from the English version.
B.3 MedMCQA
MedMCQA Pal et al. (2022) comprises a large number of questions from the Indian medical entrance exam, covering 2400 healthcare topics and 21 medical subjects. For the MIRAGE benchmark, we utilized the “dev” set of the original MedMCQA dataset.
B.4 PubMedQA*
PubMedQA is a research QA dataset in the biomedical field, consisting of 1000 manually annotated questions constructed from PubMed abstracts. In the MIRAGE benchmark, the reference contexts were removed. We selected a subset of 500 questions, which we refer to as PubMedQA*.
B.5 BioASQ-Y/N
BioASQ Tsatsaronis et al. (2015) is an annual biomedical QA competition. For the MIRAGE benchmark, we selected only the Machine Reading Comprehensive Track (Task B), focusing on 618 questions from recent years (2019-2023).
The important statistics of the copora and the QA datasets are presente in Table 3 below.
Corpus #Doc #Snip. $\bar{L}$. Domain
PubMed 23.9M 23.9M 196 Bio.-Med.
StatPearls 9.3k 301.2k 119 Clinics
Textbooks 18 125.8k 182 Medicine
Wikipedia 6.5M 29.9M 162 General
MedCorp 30.4M 54.2M 221 Mixed
Datasets Size #Opt. $\bar{L}$ Source
MMLU-Med 1,089 4 63 Exam.
MedQA-US 1,273 4 177 Exam.
MedMCQA 4,183 4 26 Exam.
PubMedQA* 500 3 24 Literature
BioASQ-Y/N 618 2 17 Literature
Table 3: Important statistics of corpora and QA datasets
Appendix C Exact Versions of LLMs
The exact versions of the backbone LLMs tested are listed in Table 4.
LLM name LLM version LLM site
ChatGPT gpt-3.5-turbo-16k https://platform.openai.com/docs/models/gpt-3.5-turbo
Llama3 Meta-Llama-3-8B https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
InternLM2 internlm2-123b https://www.sensetime.com/en/news-detail/51167237
ChatGLM3 chatglm3-6b https://huggingface.co/THUDM/chatglm3-6b
Qwen1.5 Qwen1.5-MoE-A2.7B https://huggingface.co/Qwen/Qwen1.5-MoE-A2.7B-Chat
Table 4: Versions of the backbone LLMs
Appendix D Qualitative Results of the Router
As illustrated in Figure 4, the router effectively assigns different weights to various granularity levels. From the figure, we can infer that a potential global peak in weights may exist at a granularity level smaller than level 1 or larger than level 5.
However, we did not test these smaller or larger granularity levels due to the following reasons: In our experiments, we focus exclusively on LLMs with a parameter size of around 7 billion. For these models, the chunking size at granularity level 5 approaches their maximum context window. Conversely, the chunking size at granularity level 1 consists of only a few dozen characters, which is already quite small.
Appendix E Experiment on Soft Labels
In this section, we evaluate the performance of building soft labels using different methods (TF-IDF Ramos (2003), RoBERTa Liu et al. (2019), and hitrate score). For this experiment, we fix one retriever (BM25) and the backbone LLM (Qwen1.5), then build the soft labels using these three different methods. Different routers are trained with these different soft labels and tested on various MQA datasets. The results are grouped in Table 5.
The results indicate that there is no universal best method for building soft labels. For instance, “hitrate score” is the most effective when training on MedMCQA, while RoBERTa shows advantages on PubMedQA. Overall, the soft labels built with RoBERTa demonstrate good performance across the board. Therefore, for the remainder of the experiments, RoBERTa is adopted as the default method to build soft labels.
Training datasets Methods MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
MedQA RoBERTa 0.4265$\pm$0.04 0.3711$\pm$0.04 0.3831$\pm$0.02 0.5000$\pm$0.06 0.6234$\pm$0.06 0.4608
hitrate 0.4779$\pm$0.04 0.4088$\pm$0.04 0.4330$\pm$0.02 0.4194$\pm$0.06 0.6753$\pm$0.05 0.4829
TF-IDF 0.4559$\pm$0.04 0.3019$\pm$0.04 0.4023$\pm$0.02 0.4355$\pm$0.06 0.5714$\pm$0.06 0.4334
MedMCQA RoBERTa 0.4485$\pm$0.05 0.3459$\pm$0.04 0.4023$\pm$0.02 0.4839$\pm$0.06 0.5195$\pm$0.06 0.4400
hitrate 0.4118$\pm$0.04 0.3774$\pm$0.04 0.3678$\pm$0.02 0.3387$\pm$0.06 0.5065$\pm$0.06 0.4004
TF-IDF 0.4412$\pm$0.04 0.3899$\pm$0.04 0.4119$\pm$0.02 0.5161$\pm$0.06 0.5844$\pm$0.06 0.4687
PubMedQA RoBERTa 0.4485$\pm$0.04 0.4528$\pm$0.04 0.3966$\pm$0.02 0.4677$\pm$0.06 0.6623$\pm$0.05 0.4856
hitrate 0.4118$\pm$0.04 0.4088$\pm$0.04 0.3851$\pm$0.02 0.4355$\pm$0.06 0.6104$\pm$0.06 0.4503
TF-IDF 0.4044$\pm$0.04 0.3208$\pm$0.04 0.3908$\pm$0.02 0.4677$\pm$0.06 0.5584$\pm$0.06 0.4284
Table 5: The performance of the routers trained with the soft labels created with different methods, best method marked in bold.
Appendix F Impact of the Number of Candidate Snippets
In the previous experiment, three candidate snippets were retrieved from each granularity level of the external knowledge database. In this section, we investigate the effect of varying the number of candidate snippets on the overall performance of the RAG system. The rationale behind this exploration lies in the potential limitation of a small pool of candidate snippets, such as keeping only three. In such case, valuable snippets may be overlooked. For instance, a snippet ranked fourth or fifth in each retrieval may appear repeatedly across different granularity levels and thus might be selected after its relevance scores adjusted with weights assigned by the router. However, despite potentially high relevance, such snippets are excluded early in the retrieval process.
In this experiment, we test the RAG system equipped with MoG using different numbers of candidate snippets $k_{r}$ ($k_{r}\in$3, 8, 16, 32}). For simplicity, the backbone LLM was fixed as Qwen1.5. The experiment results are shown in Figure 5.
Generally, RAG performance improves as $k_{r}$ increases, indicating the presence of helpful knowledge in the retrieved snippets. However, too many irrelevant snippets mislead the inference of LLM based on its knowledge. MoG stands out with its high initial performance even at low $k_{r}$ values ($k_{r}$<=8 in this case), thanks to its multi-granularity filtering, which efficiently selects relevant snippets. This enhances the LLM’s accuracy without requiring a large snippet pool. MoG even strikes a balance at high $k_{r}$ values ($k_{r}$>=16 in this case) due to its threshold limiting of the different corpora, avoiding the pitfalls of excessive information retrieval. MoG’s consistently high performance demonstrates its superiority in optimizing the number of candidate snippets for effective medical question-answering.
Figure 5: Accuracy of Medical Question-Answering task with different number of candidate snippets
Appendix G Choice of Retriever
In the previous experiment, BM25 Robertson and Zaragoza (2009) was used as the retriever because it is a lightweight and popular choice in practice. In this section, we replace BM25 with other popular retrievers to evaluate their performance. Referring to the setup in MedRAG Xiong et al. (2024), we select a general-domain semantic retriever called Contriever Izacard et al. (2022a), a scientific-domain retriever called SPECTER Cohan et al. (2020), and a biomedical-domain retriever called MedCPT Jin et al. (2023). Additionally, the Reciprocal Rank Fusion (RRF) method Cormack et al. (2009) is utilized to combine results from different retrievers, including RRF-2 (fusion of BM25 and MedCPT) and RRF-4 (fusion of all four retrievers). From Table 6, we can observe the following: Each retriever has a well-performing QA set, which results in minimal overall differences among these retrievers. Considering these conclusions, we decide to continue using BM25 for all other experiments.
Corpus Retriever MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
Textbooks BM25 0.6139$\pm$0.04 0.3961$\pm$0.04 0.3907$\pm$0.02 0.3100$\pm$0.06 0.5000$\pm$0.05 0.4421
Contriever 0.5644$\pm$0.03 0.4235$\pm$0.03 0.4074$\pm$0.01 0.2500$\pm$0.04 0.5161$\pm$0.04 0.4323
SPECTER 0.5941$\pm$0.03 0.4118$\pm$0.03 0.3919$\pm$0.01 0.1900$\pm$0.04 0.4839$\pm$0.04 0.4143
MedCPT 0.5347$\pm$0.04 0.4549$\pm$0.03 0.4026$\pm$0.02 0.3100$\pm$0.05 0.5565$\pm$0.04 0.4517
RRF-2 0.5396$\pm$0.03 0.4196$\pm$0.03 0.4241$\pm$0.02 0.2800$\pm$0.04 0.5403$\pm$0.04 0.4407
RRF-4 0.5495$\pm$0.04 0.4392$\pm$0.03 0.4397$\pm$0.02 0.2600$\pm$0.04 0.6129$\pm$0.04 0.4603
Table 6: Performance of different retrievers
Appendix H Performance of MoG on Medical QA Task (Other results)
In this section, we present the experiment result of MoG on the Medical QA task with four different corpora as the training datasets for the router. The results are grouped in Table 7.
LLM Method Training Corpora (Avg Acc.)
Textbooks StatPearls PubMed Wikipedia
GLM3 CoT 0.3908 0.3908 0.3908 0.3908
MedRAG 0.4269 0.4343 0.4716 0.4582
MoG 0.4492 0.4465 0.4911 0.4241
GPT-3.5 CoT 0.6571 0.6571 0.6571 0.6571
MedRAG 0.5250 0.5229 0.6322 0.5332
MoG 0.6122 0.5921 0.6795 0.6154
InternLM CoT 0.5914 0.5914 0.5914 0.5914
MedRAG 0.4223 0.4380 0.5559 0.4506
MoG 0.5800 0.5886 0.6394 0.5810
Llama3 CoT 0.6386 0.6386 0.6386 0.6386
MedRAG 0.5127 0.5133 0.6170 0.5206
MoG 0.6221 0.6415 0.6394 0.6328
Qwen1.5 CoT 0.3845 0.3845 0.3845 0.3845
MedRAG 0.4399 0.4623 0.4499 0.4391
MoG 0.6129 0.4622 0.5145 0.4547
Table 7: Accuracy of Medical Question-Answering task with MoG (trained with different corpora)
Appendix I Analysis on Samples Improved or Degraded
We counted the number of samples of four categories before and after applying MoG. The four categories are 1. improved (CoT gives wrong answer, MoG gives correct answer); 2. degraded (CoT gives correct answer, MoG gives wrong answer); 3. remain_correct (both CoT and MoG give correct answers); 4. remain_wrong (both CoT and MoG give wrong answers). The results are visualized as Figure 6. In this presented figure, the MoG is trained on the Wikipedia corpus. The pattern observed from this figure is similar to the ones shown in other test results. The zeroes in bars named “cot_err” and “mog_err” indicate that all the responses are correctly parsed when counting.
Around 10% of the degraded samples are randomly chosen and manually verified. We confirmed that in most (95%) degraded cases being verified, all the following statements are true:
  • several candidate snippets are retrieved correctly;
  • top-2 candidate snippets are correctly selected with the weights calculated by the router;
  • the prompt is correctly augmented and passed to backbone LLM;
  • the LLM generates the response without error;
  • the final choice (A/B/C/D or Yes/No) was correctly parsed.
In other words, the LLM changed its answer based on the snippet retrieved, it is infected by the introduced noises. The degradation is caused by the fact that the tested RAG system lacks a noise filtering mechanism, rather than by the default of MoG or MoGG.
Figure 6: Number of samples improved or degraded after application of MoG
Appendix J Performance of MoGG on Medical QA Task (Other results)
In this section, we present the experiment result of MoGG on the Medical QA task with StatPearls as the training corpus for the router. The results are grouped in Table 8.
LLM Method MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
GLM3 CoT 0.4901$\pm$0.04 0.3294$\pm$0.03 0.3799$\pm$0.02 0.3900$\pm$0.05 0.5645$\pm$0.04 0.4181
MedRAG 0.4901$\pm$0.04 0.3294$\pm$0.03 0.3799$\pm$0.02 0.4400$\pm$0.05 0.5323$\pm$0.04 0.4343
MoG 0.4851$\pm$0.04 0.2902$\pm$0.03 0.3799$\pm$0.02 0.4400$\pm$0.05 0.6371$\pm$0.04 0.4465
MoGG 0.4802$\pm$0.04 0.3059$\pm$0.03 0.3668$\pm$0.02 0.4600$\pm$0.05 0.5806$\pm$0.04 0.4387
GPT-3.5 CoT 0.7624$\pm$0.03 0.6706$\pm$0.03 0.5866$\pm$0.02 0.4200$\pm$0.05 0.7339$\pm$0.04 0.6347
MedRAG 0.6881$\pm$0.03 0.6549$\pm$0.03 0.4552$\pm$0.02 0.2600$\pm$0.04 0.5565$\pm$0.04 0.5229
MoG 0.7277$\pm$0.03 0.6471$\pm$0.03 0.5400$\pm$0.02 0.3200$\pm$0.05 0.7258$\pm$0.04 0.5921
MoGG 0.7624$\pm$0.03 0.6314$\pm$0.03 0.5460$\pm$0.02 0.3700$\pm$0.05 0.7177$\pm$0.04 0.6055
InternLM CoT 0.7228$\pm$0.03 0.6000$\pm$0.03 0.5352$\pm$0.02 0.3700$\pm$0.05 0.7339$\pm$0.04 0.5924
MedRAG 0.6584$\pm$0.03 0.5137$\pm$0.03 0.3596$\pm$0.02 0.1500$\pm$0.04 0.5081$\pm$0.04 0.4380
MoG 0.6881$\pm$0.03 0.5961$\pm$0.03 0.5448$\pm$0.02 0.3800$\pm$0.05 0.7339$\pm$0.04 0.5886
MoGG 0.6782$\pm$0.03 0.5725$\pm$0.03 0.3500$\pm$0.02 0.7500$\pm$0.05 0.7500$\pm$0.04 0.6201
Llama3 CoT 0.7277$\pm$0.03 0.6392$\pm$0.03 0.5663$\pm$0.02 0.5900$\pm$0.05 0.7177$\pm$0.04 0.6482
MedRAG 0.6089$\pm$0.03 0.5882$\pm$0.03 0.4170$\pm$0.02 0.3800$\pm$0.05 0.5242$\pm$0.04 0.5133
MoG 0.7376$\pm$0.03 0.5961$\pm$0.03 0.5317$\pm$0.02 0.6000$\pm$0.05 0.7419$\pm$0.04 0.6415
MoGG 0.7030$\pm$0.03 0.6275$\pm$0.03 0.5436$\pm$0.02 0.5900$\pm$0.05 0.7258$\pm$0.04 0.6380
Qwen1.5 CoT 0.4604$\pm$0.04 0.3255$\pm$0.03 0.3883$\pm$0.02 0.2000$\pm$0.04 0.5484$\pm$0.04 0.3845
MedRAG 0.5495$\pm$0.04 0.4471$\pm$0.03 0.4146$\pm$0.02 0.3600$\pm$0.05 0.5403$\pm$0.04 0.4623
MoG 0.5446$\pm$0.04 0.3882$\pm$0.03 0.4337$\pm$0.02 0.3400$\pm$0.05 0.6048$\pm$0.04 0.4623
MoGG 0.5446$\pm$0.04 0.4157$\pm$0.03 0.4253$\pm$0.02 0.3500$\pm$0.05 0.6290$\pm$0.04 0.4729
Table 8: Accuracy of Medical Question-Answering task with MoGG (trained with StatPearls), best results marked in bold
Appendix K MoGG’s Abaltion test
In this section, we rerun the MoGG without the router on Llama 3 to highlight its effectiveness. The results of MoGG’s ablation test is in Table 9 (highlighted in bold). The results are similar to MedRAG, because, without the router, MoGG is essentially the same as MedRAG, the added flexibility of the graph will only show when used with a router.
Method MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
CoT 0.7079 0.6431 0.5663 0.5500 0.7258 0.6386
MedRAG 0.6485 0.5961 0.4146 0.3800 0.5242 0.5127
MoG 0.7228 0.6196 0.5484 0.5100 0.7097 0.6221
MoGG w/o router 0.6683 0.5451 0.4182 0.3800 0.5484 0.5120
MoGG 0.7070 0.5961 0.5460 0.5200 0.7661 0.6262
Table 9: Ablation test with Llama3 as backbone LLM
Appendix L MoG Trained With Only Textbooks Corpus
In this Appendix, we present the results of training MoG only with Textbooks corpus. As Textbooks are only a subset of MedCorp, the performance is not as good as the ones shown in Table 1. Compared with Table 10 and Table 2, we can see the advantage of MoGG more clearly.
LLM Method MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
GLM3 MedRAG 0.4802 0.3569 0.3811 0.3600 0.5565 0.4269
MoG 0.5545 0.2941 0.3548 0.4700 0.5726 0.4492
GPT-3.5 MedRAG 0.7277 0.6745 0.4468 0.2600 0.5161 0.5250
MoG 0.7525 0.6667 0.5603 0.5200 0.7016 0.6122
InternLM MedRAG 0.6188 0.5569 0.3118 0.1400 0.4839 0.4223
MoG 0.7277 0.5725 0.5281 0.3400 0.7339 0.5804
Llama3 MedRAG 0.6485 0.5961 0.4146 0.3800 0.5242 0.5127
MoG 0.7228 0.6196 0.5484 0.5100 0.7097 0.6221
Qwen1.5 MedRAG 0.5941 0.4000 0.3835 0.3300 0.4919 0.4399
MoG 0.5792 0.3843 0.4110 0.3300 0.6129 0.4635
Table 10: Accuracy of Medical Question-Answering task with MoG (trained with only Textbooks)
Appendix M Detailed Analysis of Time Consumption
In this appendix, we present the detailed experiment results of execution time. We measure the wall-clock execution time with different number of granularities using Llama3 as the backbone LLM. The results organized in Table 11 show that:
(1) the introduction of the router module will effectively increase the inference time by about 60%;
(2) increasing the number of granularity levels will only increase a marginal increase in execution time. The reason is that the bottleneck of execution time is the API calling time of backbone LLMs.
# (Granularity) MMLU MedQA MedMCQA PubMedQA BioASQ Global Avg.
1 w/o router 6.61 9.02 10.01 7.40 7.53 8.12
1 w/ router 10.15 10.76 15.81 15.99 13.09 13.16
2 13.25 12.60 20.53 17.30 14.98 15.73
3 16.30 16.01 19.03 16.73 17.12 17.04
4 13.43 20.65 22.72 18.93 14.05 17.96
5 13.20 20.43 23.00 19.36 13.66 17.93
Table 11: Performance Metrics by Granularity